OpenStudy (anonymous):

Write the expression for the slope of a line whose equation is Ax+By+C=0?

OpenStudy (akashdeepdeb):

Slope is normally given in the equation y = mx + b [WHERE m is slope okay?] [AND y CO-EFF. IS ZERO] So in this equation Ax + By + c = 0 or By = - Ax - c And then divide the LHS n RHS by B So what do YOU think m or the slope will be here? Refer to the equation y = mx+ b! Try! :)

OpenStudy (anonymous):

so its y=(-Ax-c/B) ?

OpenStudy (akashdeepdeb):

Yeah kinda! y = -Ax/B - c/B So in the equation y = mx + c m is the co-efficient of x and the slope! What do YOU think the slope will be here? :D

OpenStudy (akashdeepdeb):

@Ailee

OpenStudy (anonymous):

o-o errr I thought it would just be -Ax/b. sorry im confused D:

OpenStudy (akashdeepdeb):

CHILLOUT!! :D Good try though! [I wasn't have as good as you when I WAS LEARNING THIS] -Ax/B is the whole thing [-A/B is the coefficient] That means that -A/B is the constant term getting me? So here the slop of the equation Ax + By + c = 0 Should be?? m = -A/B Because it is the co-efficient of x written in the form = y = -Ax/B - c/B Understood?? :D

OpenStudy (anonymous):

YES!!:D i get that part but then what to do after that step?

OpenStudy (akashdeepdeb):

Say we have the equation 2x - 3y = 6 okay? Try finnding the slope (m) of that!

OpenStudy (anonymous):

the slope is 2/3x

OpenStudy (akashdeepdeb):

Good Good! But you see there is only one problem the slope is always the constant term Not including the variable so the slop would be ? 2/3!! :D [YIPPEEE] Did you get this question now? :)

OpenStudy (anonymous):

OOHHH SO if im not mistaken it should be y=-A/B - C/B ??? O-O

OpenStudy (akashdeepdeb):

No Ailee [LISTEN TO ME CAREFULLY] See y = mx + b --> IS THE STANDARD FORM OF 'ANY' LINE [Where m is slope] So in our case By = -Ax - c or y = -Ax/B - c/B \[y = mx + b\] In our case though \[y = -Ax/B - c/B\] So in our case our slope will be? -A/B because it is like the 'm' in the standard line equation! Getting me? Ask me if you have any question. :D

OpenStudy (anonymous):

so the answer is -A/B because its the slope in the equation ?

OpenStudy (anonymous):

and they are also only looking for the slope

OpenStudy (akashdeepdeb):

YES!! :) YOU ARE RIGHT!! :D

OpenStudy (anonymous):

OMG YAY!!! :DDDDDDDDDDDDDDDDDDD

OpenStudy (anonymous):

YOUR SO COOL :3

OpenStudy (akashdeepdeb):

Thank you! :D :) Have fun!! :D

OpenStudy (anonymous):

If you could help, I have another question I don't know how to start: write the standard form form of the equation of the line that is parallel to the graph of the given equation and passes through the point with the given 4x-9y= -23;(12,-15)

OpenStudy (akashdeepdeb):

Hold on for one sec!

OpenStudy (akashdeepdeb):

Ok now see 

OpenStudy (akashdeepdeb):

Both the lines obviously have same gradient only and only because they are parrllel. [FIND THE SLOPE LIKE I SHOWED YOU] Thus, if they are parrallel slope of the parrallel line passes through (12,-15) and has slope 4/9 Now there is something known as the point slope equation of the line! That states: (y-y1) = m(x-x1) So we have (x1,y1) = (12,-15) And m = 4/9 Put it in the line equation and then get the equation of the parallel line!! :D Understood?? :)

OpenStudy (anonymous):

ok, thanks C;

OpenStudy (akashdeepdeb):

Medal me plz?

OpenStudy (akashdeepdeb):

:/